OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2.

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

y=mx+b is the eqaution of a line so now we sub in the values -2=2(3)+b b=-8 so the equation of the line is y=2x-8